A 155 g ball is tied to a string. It is pulled to an angle of 7.80 ∘ and released to swing as a pendulum. A student with a stopwatch finds that 10 oscillations take 11.0 s . How long is the string?
Answer #1

here,

time period T = 2 *pi sqrt[L/g]

this relation is valid for small angles of oscillations (7.80 < 10 degree) qualifies that

10 oscillations take time = 11 sec

1 oscillation will take time = 11/10 sec = T

T = 1.1 (sec)

L = length string = g *[T/2pi]^2

L = 9.8 * [0.1 * 3.14]^2

L = 0.97 meter

length of string is 0.97 m